Find the limit of the sequence section 3.2.6

1
Expert's answer
2022-01-24T16:17:12-0500

"\displaystyle
\text{Consider the sequence, $\frac{n-1}{n}$}\\
\lim_{n \to \infty}\frac{n-1}{n} = \lim_{n \to \infty} (1-\frac1n)\\
=\lim_{n \to \infty}1 - \lim_{n \to \infty} \frac1n\\
\text{Where $\lim_{n \to \infty} \frac1n = 0$}\\
\text{Hence,} \lim_{n \to \infty}\frac{n-1}{n}=1"